The slope of the line passing through points (5,0) and (7,4)
Answer by mananth(16949)

The slope of the line passing through points (5,0) and (7,4)
slope = m = (y2-y1)/(x2-x1) = (4-0)/7-5)
=4/2
= 2
slope = 2